An L-R-C series circuit L = 0.121 H , R = 235 Ω , and C = 7.34 μF carries an rms current of 0.455 A with a frequency of 398 Hz .

a) What average power is delivered by the source?

b) What is the average rate at which electrical energy is converted to thermal energy in the resistor?

c) What is the average rate at which electrical energy is dissipated (converted to other forms) in the capacitor?

d) What is the average rate at which electrical energy is dissipated (converted to other forms) in the inductor?
Answer #1

given

L= 0.121 H

f= 398Hz

X_{L}=wL

X_{L}=2\pi fL

X_{L}=2\pi \times 398\times 0.121\Omega

X_{L}=302.59\Omega

C=7.34*10-6 F

X_{C}=\frac{1}{2\pi fC}\Omega

X_{C}=\frac{1}{2\pi\times 398\times 7.34\times 10^{-6}}\Omega

X_{C}=54.480\Omega

so the impedance

Z=\sqrt{R^{2}+(X_{L}-X_{C})^{2}}

Z=\sqrt{235^{2}+(302.59-54.480)^{2}}

Z=341.73\Omega

voltage

V=IZ

a)

power is

P=IV

P=I^{2}Z

P=0.455^{2}\times 341.73

P=70.7W

++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++

b)

for the resistor

VR= IR

so the power that converted to electrical energy

P_{R}=I^{2}R

P_{R}=0.455^{2}\times 398W

P_{R}=82.4W

++++++++++++++++++++++++++++++

for the capacitor

P_{C}=I^{2}X_{C}

P_{C}=0.455^{2}\times 54.480

P_{C}=11.3W

+++++++++++++++++++++++++++++++++++++

power in inductor

P_{L}=I^{2}L

P_{L}=0.455^{2}\times 302.59W

P_{L}=62.6W